OpenStudy (anonymous):

Divide: 7a + 28/ a-4 divided by a^2 + a -12/ 14a

OpenStudy (anonymous):

\[\frac{(7a+28)(14a)}{(a-4)(a^2+a-12)}\] is a start

OpenStudy (anonymous):

then factor and cancel if you can

OpenStudy (anonymous):

i need help factoring

OpenStudy (anonymous):

\(7a+28=7(a+4)\) for the numerator

OpenStudy (anonymous):

\(a^2+a-12=(a+4)(a-3)\) for the denominator

OpenStudy (anonymous):

what would be the next step?

OpenStudy (anonymous):

@satellite73